Circular No. 7484

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 Q1 (SOHO) Further to IAUC 7445 and 7480, D. Biesecker reports positions by

D. Hammer for a comet found by J. Danaher in LASCO C3 images at the SOHO website. No tail is evident, and Biesecker provides some V magnitude estimates (in addition to the discovery magnitude given below): Aug. 28.638 UT, 8.0; 28.846, 7.2; 29.071, 7.7; 29.279, 8.4; 29.513, 7.9; 29.779, 10.3. The reduced measurements and parabolic orbit by B. G. Marsden are given on MPEC 2000-Q42.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 28.473  10^{h}48.^{m}1  + 8^{o}38′  8.6

COMET C/2000 OF_{8} (SPACEWATCH) An object reported by Spacewatch as asteroidal (MPEC 2000-P03;

discovery observation below) appears slightly diffuse in individual CCD images taken by O. R. Hainaut and C. E. Delahodde with the ESO/MPI 2.2-m telescope on Aug. 29.00 UT (3′′-4′′ coma extending northeastward); the coma is clearly visible on a 30-min stacked image. D. D. Balam also now finds a 6′′ coma elongated in p.a. 260^{o} in stacked images taken on Aug. 28.27 with the 1.82-m Plaskett reflector. See also MPEC 2000-Q43.

2000 UT  α_{2000}  δ_{2000}  m_{2}
July 24.32235  20^{h}57^{m}31.^{s}48  −7^{o}08′18′′.3  21.3

SUPERNOVA 2000dg IN MCG +1-1-29 S. Benetti, A. Zacchei, and I. Perez, Telescopio Nazionale Galileo

(TNG), report: “A fully reduced CCD spectrum (range 382.4–811.1 nm, resolution 1.3 nm), obtained on Aug. 28.13 UT with the TNG reflector (+ D.o.lo.res spectrograph), shows SN 2000dg (cf. IAUC 7480) to be a type-Ia supernova at maximum light. The spectrum consists of a blue continuum with superimposed P-Cyg lines of Ca II (H and K), S II, and Si II ions. The expansion velocity deduced from minimum of the Si II 635.5-nm doublet is 10 850 km/s (a recession velocity for the parent galaxy of 11 542 km/s was obtained from the LEDA database).”

A. V. Filippenko and W. D. Li, University of California, Berkeley, report that inspection of a CCD spectrum (range 330–1000 nm), obtained on Aug. 27 UT with the Shane 3-m reflector at Lick Observatory, also shows SN 2000dg to be of type Ia, near maximum brightness.

2000 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green